# Measure theoretic properties of large products of consecutive partial quotients

### v1

@@ 4.2. Divergence case
$\mathfrak{m}(U_{1}(m,n))$ $\leq m\frac{\log^{\ell-1}\varphi(m)}{\varphi(m)}\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\ell\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}$
$\ll\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}.$
This shows (18). Note that we have used that for the values of $k$ considered the blocks $(a_{k},\ldots,a_{k+\ell-1})$ and $(a_{n},\ldots,a_{n+\ell-1})$ do not overlap. Here and in the following, whenever we assert something about blocks of a sequence of natural numbers without quantifiers, it should be understood that the assertion holds for every sequence of natural numbers.
We consider two different series to estimate $\mathfrak{m}(U_{2}(m,n))$. First, if $M\leq j\leq m-\ell$ then the blocks $(a_{j},\ldots,a_{j+\ell-1})$ and $(a_{m},\ldots,a_{m+\ell-1})$ do not overlap and
$\sum_{j=M}^{m-\ell}\sum_{\begin{subarray}{c}a_{j}\cdots a_{j+\ell-1}\geq\varphi(m)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{j+\ell-1}^{2}a_{m}^{2}\cdots a_{m+\ell-1}^{2}}$ $\ll m\frac{\log^{\ell-1}\varphi(m)}{\varphi(m)}\frac{\log^{\ell-1}\varphi(n)}{\varphi(n)}$
$\leq\frac{\log^{\ell-1}\varphi(n)}{\varphi(n)}.$
The product $a_{j}^{2}\cdots a_{m+\ell-1}^{2}$ in the series to come should be interpreted as follows: if $j+\ell-1\leq m$, then $a_{j}^{2}\cdots a_{m+\ell-1}^{2}$ is the product of the products of the blocks $(a_{j},\ldots,a_{j+\ell-1})$ and $(a_{m},\ldots,a_{m+\ell-1})$; if $j+\ell-1>m$, then $a_{j}^{2}\cdots a_{m+\ell-1}^{2}$ is the product of the block $(a_{j},a_{j+1},\ldots,a_{m+\ell-1})$. We use a similar convention for other combinations of the parameters $j,k,m,n$.
When $m-\ell+1\leq j\leq m-1$, the blocks $(a_{j},\ldots,a_{j+\ell-1})$ and $(a_{m},\ldots,a_{m+\ell-1})$ overlap and
$\sum_{j=m-\ell+1}^{m-1}\sum_{\begin{subarray}{c}a_{j}\cdots a_{j+\ell-1}\geq\varphi(m)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+\ell-1}^{2}}$ $\leq\sum_{j=m-\ell+1}^{m-1}\sum_{a_{m}\cdots a_{m+\ell-1}\geq\varphi(n)}\frac{1}{a_{j}^{2}\cdots a_{m+\ell-1}^{2}}$
$\ll_{\ell}\frac{\log^{\ell-1}\varphi(n)}{\varphi(n)}.$
We may now argue as in the convergence case recalling the restriction $a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)$ to conclude (19).
Next, we show (4.2). First, we consider the values of $k$ for which the block $(a_{k},\ldots,a_{k+\ell-1})$ does not overlap neither $(a_{n},\ldots,a_{n+\ell-1})$ nor $(a_{m},\ldots,a_{m+\ell-1})$. In other words, we assume $m+\ell\leq k\leq n-\ell$. This leads us to the series
$\sum_{j=M}^{m-1}\sum_{k=m+\ell}^{n-\ell}\sum_{\begin{subarray}{c}a_{j}\cdots a_{j+\ell-1}\geq\varphi(m)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(m)\\ a_{k}\cdots a_{k+\ell-1}\geq\varphi(n)\\ a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+\ell-1}^{2}a_{k}^{2}\cdots a_{k+\ell-1}^{2}a_{n}^{2}\cdots a_{n+\ell-1}^{2}},$
@@
$\left(\frac{\log^{\ell-1}\varphi(n)}{\varphi(n)}\right)^{2}$ $\sum_{j=M}^{m-1}\sum_{\begin{subarray}{c}a_{j}\cdots a_{j+\ell-1}\geq\varphi(m)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(m)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+\ell-1}^{2}}\ll$
$\ll\left(\frac{\log^{\ell-1}\varphi(n)}{\varphi(n)}\right)^{2}\left(m\frac{\log^{2(\ell-1)}\varphi(m)}{\varphi(m)}+\frac{\log^{\ell-2}\varphi(m)}{\varphi(m)}\right).$
The last estimate is obtained as in the convergence case. Now suppose that $n-\ell+1\leq k\leq n-1$. This means that $(a_{k},\ldots,a_{k+\ell-1})$ and $(a_{n},\ldots,a_{n+\ell-1})$ overlap but $(a_{k},\ldots,a_{k+\ell-1})$ and $(a_{m},\ldots,a_{m+\ell-1})$ do not (by the definition of $\mathcal{I}_{1}$), hence
$\sum_{j=M}^{m-1}$ $\sum_{k=n-\ell+1}^{n-1}\sum_{\begin{subarray}{c}a_{j}\cdots a_{j+\ell-1}\geq\varphi(m)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(m)\\ a_{k}\cdots a_{k+\ell-1}\geq\varphi(n)\\ a_{n}\cdots a_{n+\ell-1}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+\ell-1}^{2}a_{k}^{2}\cdots a_{n+\ell-1}^{2}}\ll$
$\ll\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}\sum_{j=M}^{m-1}\sum_{\begin{subarray}{c}a_{j}\cdots a_{j+\ell-1}\geq\varphi(m)\\ a_{m}\cdots a_{m+\ell-1}\geq\varphi(m)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+\ell-1}^{2}}$
$\ll\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}\left(m\frac{\log^{2(\ell-1)}\varphi(m)}{\varphi(m)}+\frac{\log^{\ell-2}\varphi(m)}{\varphi(m)}\right).$
Lastly, we deal with the values of $k$ for which the blocks $(a_{m},\ldots,a_{m+\ell-1})$ and $(a_{k},\ldots,a_{k+\ell-1})$ overlap. This happens if and only if $k\leq m+\ell-1$. In this case, $k+\ell-1\leq m+2(\ell-1)<n$, hence $(a_{k},\ldots,a_{k+\ell-1})$ and $(a_{n},\ldots,a_{n+\ell-1})$ do not overlap. For any $j\in\{M,\ldots,m-1\}$, $k\in\{m+1,\ldots,m+\ell-1\}$ define
$D_{j,k}:=\left\{x\in I_{t}:\begin{matrix}a_{j}(x)\cdots a_{j+\ell-1}(x)\geq\varphi(m),&a_{m}(x)\cdots a_{m+\ell-1}(x)\geq\varphi(m),\\ a_{k}(x)\cdots a_{k+\ell-1}(x)\geq\varphi(n),&a_{n}(x)\cdots a_{n+\ell-1}(x)\geq\varphi(n)\end{matrix}\right\}.$
@@
$\bigcup_{j=M}^{m-1}D_{j,k}\subseteq\left\{x\in I_{t}:\begin{matrix}a_{k}(x)\cdots a_{k+\ell-1}(x)\geq\varphi(n),\\ a_{n}(x)\cdots a_{n+\ell-1}(x)\geq\varphi(n)\end{matrix}\right\}=A_{n,k}\cap I_{t}.$
@@
$\mathfrak{m}\left(\bigcup_{k=m+1}^{m+\ell-1}\bigcup_{j=M}^{m-1}D_{j,k}\right)\leq\sum_{k=m+1}^{m+\ell-1}\mathfrak{m}(A_{n,k}\cap I_{t})\ll_{\ell}\frac{\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}\mathfrak{m}(I_{t})$
@@
$\mathfrak{m}\left(\limsup_{n\to\infty}E_{n}\right)\gg_{\ell}\mathfrak{m}(I_{t}).$
@@
The cases $B=1$ and $B=\infty$ follow from $\mathcal{F}_{2}(\varphi)\subseteq\mathcal{F}_{3}(\varphi)\subseteq\mathcal{E}_{2}(\varphi)$ and theorems 1.8 and 1.2. Thus, we assume $1<B<\infty$. Define the functions $X_{1},X_{2},X_{3}:[0,1]\to\mathbb{R}$ by
@@ 5.1.1. Overlaps of length 2
We cover $\mathcal{F}_{3,n}^{2}(B)$ by the sets $\mathcal{F}_{3,n}^{1,C}(B)$ and $\mathcal{F}_{3,n}^{1,G}(B)$ given by
@@ 6. Final remarks and further results
With regards to the Hausdorff dimension of certain Lebesgue null sets, for an increasing function $\phi:\mathbb{N}\to\mathbb{R}$ with $\phi(n)\to\infty$ as $n\to\infty$ we consider the set
@@
Recall that the $n$-th term of the series in Theorem 1.5 is
$\frac{n\log^{2(\ell-1)}\varphi(n)}{\varphi^{2}(n)}+\frac{\log^{\ell-2}\varphi(n)}{\varphi(n)}.$
It is clear from the proof that the first term accounts for the large products appearing in non-overlapping blocks and the second term corresponds overlapping blocks. Hence, for $\ell=1$ the second term is non-existent.

### v2

@@ 4.2. Divergence case
$\mathfrak{m}(U_{1}(m,n))$ $\leq m\frac{\log^{2}\varphi(m)}{\varphi(m)}\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}+3\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}$
$\ll\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}.$
This shows (20). Note that we have used that for the values of $k$ considered the blocks $(a_{k},a_{k+1},a_{k+2})$ and $(a_{n},a_{n+1},a_{n+2})$ do not overlap. Here and in the following, whenever we assert something about blocks of a sequence of natural numbers without quantifiers, it should be understood that the assertion holds for every sequence of natural numbers.
We consider two different series to estimate $\mathfrak{m}(U_{2}(m,n))$. First, if $M\leq j\leq m-3$ then the blocks $(a_{j},a_{j+1},a_{j+2})$ and $(a_{m},a_{m+1},a_{m+2})$ do not overlap and
$\sum_{j=M}^{m-\ell}\sum_{\begin{subarray}{c}a_{j}a_{j+1}a_{j+2}\geq\varphi(m)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}a_{j+1}^{2}a_{j+2}^{2}a_{m}^{2}a_{m+1}^{2}a_{m+2}^{2}}$ $\ll m\frac{\log^{2}\varphi(m)}{\varphi(m)}\frac{\log^{2}\varphi(n)}{\varphi(n)}$
$\leq\frac{\log^{2}\varphi(n)}{\varphi(n)}.$
The product $a_{j}^{2}\cdots a_{m+2}^{2}$ in the series to come should be interpreted as follows: if $j+2\leq m$, then $a_{j}^{2}\cdots a_{m+2}^{2}$ is the product of the products of the blocks $(a_{j},a_{j+1},a_{j+2})$ and $(a_{m},a_{m+1},a_{m+2})$; if $j+2>m$, then $a_{j}^{2}\cdots a_{m+2}^{2}$ is the product of the block $(a_{j},a_{j+1},\ldots,a_{m+2})$. We use a similar convention for other combinations of the parameters $j,k,m,n$.
When $m-2\leq j\leq m-1$, the blocks $(a_{j},a_{j+1},a_{j+2})$ and $(a_{m},a_{m+1},a_{m+2})$ overlap and
$\sum_{j=m-\ell+1}^{m-1}\sum_{\begin{subarray}{c}a_{j}a_{j+1}a_{j+2}\geq\varphi(m)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+2}^{2}}$ $\leq\sum_{j=m-2}^{m-1}\sum_{a_{m}a_{m+1}a_{m+2}\geq\varphi(n)}\frac{1}{a_{j}^{2}\cdots a_{m+2}^{2}}$
$\ll\frac{\log^{2}\varphi(n)}{\varphi(n)}.$
We may now argue as in the convergence case recalling the restriction $a_{n}a_{n+1}a_{n+2}\geq\varphi(n)$ to conclude (21).
Next, we show (4.2). First, we consider the values of $k$ for which the block $(a_{k},a_{k+1},a_{k+2})$ does not overlap neither $(a_{n},a_{n+1},a_{n+2})$ nor $(a_{m},a_{m+1},a_{m+2})$. In other words, we assume $m+3\leq k\leq n-3$. This leads us to the series
$\sum_{j=M}^{m-1}\sum_{k=m+3}^{n-3}\sum_{\begin{subarray}{c}a_{j}a_{j+1}a_{j+2}\geq\varphi(m)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(m)\\ a_{k}a_{k+1}a_{k+2}\geq\varphi(n)\\ a_{n}a_{n+1}a_{n+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+\ell-1}^{2}a_{k}^{2}a_{k+1}^{2}a_{k+2}^{2}a_{n}^{2}a_{n+1}^{2}a_{n+2}^{2}},$
@@
$\left(\frac{\log^{2}\varphi(n)}{\varphi(n)}\right)^{2}$ $\sum_{j=M}^{m-1}\sum_{\begin{subarray}{c}a_{j}a_{j+1}a_{j+2}\geq\varphi(m)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(m)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+2}^{2}}\ll$
$\ll\left(\frac{\log^{2}\varphi(n)}{\varphi(n)}\right)^{2}\left(m\frac{\log^{4}\varphi(m)}{\varphi^{2}(m)}+\frac{\log\varphi(m)}{\varphi(m)}\right).$
The last estimate is obtained as in the convergence case.
Now suppose that $n-2\leq k\leq n-1$. This means that $(a_{k},a_{k+1},a_{k+2})$ and $(a_{n},a_{n+1},,a_{n+2})$ overlap but $(a_{k},a_{k+1},a_{k+2})$ and $(a_{m},a_{m+1},a_{m+2})$ do not (by the definition of $\mathcal{I}_{1}$), hence
$\sum_{j=M}^{m-1}$ $\sum_{k=n-2}^{n-1}\sum_{\begin{subarray}{c}a_{j}a_{j+1}a_{j+2}\geq\varphi(m)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(m)\\ a_{k}a_{k+1}a_{k+2}\geq\varphi(n)\\ a_{n}a_{n+1}a_{n+2}\geq\varphi(n)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+2}^{2}a_{k}^{2}\cdots a_{n+2}^{2}}\ll$
$\ll\frac{\log\varphi(n)}{\varphi(n)}\sum_{j=M}^{m-1}\sum_{\begin{subarray}{c}a_{j}a_{j+1}a_{j+2}\geq\varphi(m)\\ a_{m}a_{m+1}a_{m+2}\geq\varphi(m)\end{subarray}}\frac{1}{a_{j}^{2}\cdots a_{m+2}^{2}}$
$\ll\frac{\log\varphi(n)}{\varphi(n)}\left(m\frac{\log^{4}\varphi(m)}{\varphi(m)}+\frac{\log\varphi(m)}{\varphi(m)}\right).$
Lastly, we deal with the values of $k$ for which the blocks $(a_{m},a_{m+1},a_{m+2})$ and $(a_{k},a_{k+1},a_{k+2})$ overlap. This happens if and only if $k\leq m+2$. In this case, $k+2\leq m+4<n$, hence $(a_{k},a_{k+1},a_{k+2})$ and $(a_{n},a_{n+1},a_{n+2})$ do not overlap. For any $j\in\{M,\ldots,m-1\}$, $k\in\{m+1,m+2\}$ define
$D_{j,k}:=\left\{x\in I_{t}:\begin{matrix}a_{j}(x)a_{j+1}(x)a_{j+2}(x)\geq\varphi(m),&a_{m}(x)a_{m+1}(x)a_{m+2}(x)\geq\varphi(m),\\ a_{k}(x)a_{k+1}(x)a_{k+2}(x)\geq\varphi(n),&a_{n}(x)a_{n+1}(x)a_{n+2}(x)\geq\varphi(n)\end{matrix}\right\}.$
@@
$\bigcup_{j=M}^{m-1}D_{j,k}\subseteq\left\{x\in I_{t}:\begin{matrix}a_{k}(x)a_{k+1}(x)a_{k+2}(x)\geq\varphi(n),\\ a_{n}(x)a_{n+1}(x)a_{n+2}(x)\geq\varphi(n)\end{matrix}\right\}=A_{n,k}\cap I_{t}.$
@@
$\mathfrak{m}\left(\bigcup_{k=m+1}^{m+1}\bigcup_{j=M}^{m-1}D_{j,k}\right)\leq\sum_{k=m+1}^{m+1}\mathfrak{m}(A_{n,k}\cap I_{t})\ll\frac{\log^{4}\varphi(n)}{\varphi^{2}(n)}\mathfrak{m}(I_{t})$
@@
$\mathfrak{m}\left(\limsup_{n\to\infty}E_{n}\right)\gg\mathfrak{m}(I_{t}).$
@@
The cases $B=1$ and $B=\infty$ follow from $\mathcal{F}_{2}(\varphi)\subseteq\mathcal{F}_{3}(\varphi)\subseteq\mathcal{E}_{3}(\varphi)$ and theorems 1.8 and 1.2. Thus, we assume $1<B<\infty$. Define the functions $X_{1},X_{2},X_{3}:[0,1]\to\mathbb{R}$ by
@@ 5.1.1. Overlaps of length 2
We cover $\mathcal{F}_{3,n}^{2}(B)$ by the sets $\mathcal{F}_{3,n}^{2,C}(B)$ and $\mathcal{F}_{3,n}^{2,G}(B)$ given by
@@ 6. Final remarks and further results
Regarding the Hausdorff dimension of certain Lebesgue null sets, for an increasing function $\phi:\mathbb{N}\to\mathbb{R}$ with $\phi(n)\to\infty$ as $n\to\infty$ we consider the set
@@
Recently, Song, Yu, and Yu in [25] investigated the Hausdorff dimension of a related set concerning the product of two partial quotients. More precisely, for any $x\in[0,1)\setminus\mathbb{Q}$ and $n\in\mathbb{N}$ write.
$L_{n}(x)=\max_{1\leq i\leq n}\,a_{i}(x)a_{i+1}(x),$
Given a non-decreasing function $\varphi:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ satisfying certain regularity properties, they computed the Hausdorff dimension of
$L(\varphi):=\left\{x\in(0,1):\lim_{n\to\infty}\frac{L_{n}(x)}{\varphi(n)}=1\right\}.$
Their work lead us to the following problem.
Problem 1.
Let $\ell\in\mathbb{N}_{\geq 3}$. Can we extend the results in [25] to the product of $\ell\in\mathbb{N}_{\geq 3}$ consecutive partial quotients? That is, for $x\in[0,1)\setminus\mathbb{Q}$ and $n\in\mathbb{N}$, define
$L_{\ell,n}(x)=\max_{1\leq i\leq n}\,a_{i}(x)a_{i+1}(x)\cdots a_{i+\ell-1}(x).$
Given $\rho>0$ and a function $\varphi:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ such that $\log\varphi$ is regularly increasing of index $\rho$, determine the Hausdorff dimension of the set
$L_{\ell}(\varphi):=\left\{x\in(0,1):\lim_{n\to\infty}\frac{L_{\ell,n}(x)}{\varphi(n)}=1\right\}.$
@@
The $n$-th term of one of the series in Theorem 1.5 is
$\frac{n\log^{4}\varphi(n)}{\varphi^{2}(n)}+\frac{\log\varphi(n)}{\varphi(n)}.$
The first term accounts for the large products appearing in non-overlapping blocks and the second term corresponds overlapping blocks. Hence, for $\ell=1$ the second term is non-existent.